# Probabilistic Predictions of Process-Induced Deformation in Carbon/Epoxy Composites Using a Deep Operator Network

### v2

@@ 1 Introduction
Thermoset fiber reinforced composites are widely adopted for lightweight structural applications in aerospace[1_Sudhin2020], wind energy [2_murray2021], automotive[3_Deshpande2022, 4_deshpande2023, 5_pradeep2024], prosthetics, wearable electronics and sports equipment. A common route for their manufacturing is the lay-up of pre-impregnated sheets of unidirectional fibers and subsequently curing them under elevated pressure and temperature. During this process, residual stresses are induced due to the mismatch between the thermal expansion coefficients of the fiber and the matrix, as well as the shrinkage of the thermoset matrix as it cures and cross-links, thus affecting the mechanical properties of the manufactured component [6_White1993, 7_Li2014, 8_Zhao2006, 9_Agius2016, 10_Hahn1976]. Release of some of these residual stresses due to material compliance and demolding results in process induced deformation (PID) that can render the manufactured part out of design tolerances and defective. Additional stresses may also be induced if the deformed and out-of-tolerance part is assembled and fastened in place, potentially leading to premature or catastrophic failure. Accounting for PID and designing the cure cycle for PID reduction is therefore a critical aspect of the design and manufacturing of composite parts. Residual stresses have been shown to reduce by as much as 25 % - 30 % through optimization of the cure cycle [25_White1993]. Several studies have tried to develop predictive capabilities using numerical methods. Constitutive laws have been proposed and demonstrated as part of finite element schemes to model cure dependent viscosity development and evolution of modulus and corresponding residual stresses [11_Li2018, 12_Ding2016, 13_Liu2021, 14_Zobeiry2010]. White et al. [6_White1993] investigated, both experimentally and analytically (on LamCure), the effect of changing the manufacturer recommended cure cycle on the resulting curvature for asymmetric cross-ply laminates. Li et al. [17_Li2021] utilized a FE model for the multi-physics simulation to compute residual stresses, which was then modulated by an improved genetic algorithm to inversely determine the optimal process parameters to minimize residual stresses. Wang et al. [15_Wang2022] performed a fully coupled thermomechanical analysis to model the influence of cure cycles on residual stresses and degree of cure (DOC). They inferred that the temperature and duration of the second dwell had the greatest effect on residual stresses. Hui et al. [16_Hui2022] developed a collaborative multi-objective optimization strategy combining finite element-based cure process analysis with NGSA-II to demonstrate that the temperature gradient, residual stress, and process time can be reduced simultaneously.
@@
Along with advances in physics-based modeling, rapid progress in machine learning (ML) has transformed manufacturing research by enabling efficient data-driven modeling, prediction, and optimization [behbahani2022machine, ademujimi2017review, park2025machine, kiyani2023designing]. Deep learning, in particular, has become a central tool across science and engineering. A growing body of work focuses on operator learning, which aims to approximate mappings between input and output functions, a setting that naturally arises in many scientific and engineering applications. For PID in composite structures, a range of ML architectures has been explored. Lin et al. [lin2021use] used neural networks, support vector regression, and $k$-nearest neighbors models to rapidly estimate differences in DoC from process parameters, while Fan et al. [fan2023deep] employed convolutional neural networks to predict PID fields for laminates with varying stacking sequences. Theory-guided and simulation-informed approaches have also been proposed, including theory-guided ML with Gaussian process regression for process optimization [schoenholz2024accelerated], random forests for optimal cure design in co-cured composite structures [2022_Lavaggi], and simulation frameworks that capture competing mechanisms of residual stress evolution [18_Limaye2024]. More recently, Liu et al. [liu2024rapid] combined viscoelastic finite element analysis, feature selection (Random Forest, LASSO), and artificial neural networks for rapid PID prediction, while Zhang et al. [zhang2025deep] introduced deep learning models tailored to complex thermo–chemo–mechanical interactions.
@@
Operator learning provides a general ML framework for approximating mappings between function spaces rather than finite-dimensional vectors [lu2019deeponet, lu2024bridging, shih2024transformers, kovachki2023neural, CiCP-35-1194, bora2025retrofitting]. A prototypical architecture for this task is the Deep Operator Network (DeepONet) [lu2019deeponet], which employs a branch–trunk structure, the branch network encodes the input function, and the trunk network parameterizes the dependence on spatial or temporal query points. Several alternative neural operator architectures have been proposed, including Fourier Neural Operators (FNO) [li2020fourier] and U-Net-based operator models [ronneberger2015u], which mainly differ in how they represent and propagate functional information. Physics-informed variants of DeepONet incorporate PDE residuals and other physical constraints into the loss function, enabling physically consistent predictions in data-scarce regimes [wang2021learning, kiyani2025predicting]. The Two-step DeepONet [lee2024training, kiyani2025predicting, peyvan2024riemannonets] extends this idea by composing multiple operator stages, thus increasing expressivity for complex multiscale or highly nonlinear mappings while retaining an operator-centric viewpoint.
@@
Building on these developments, Feature-wise Linear Modulation (FiLM) [perez2018film] is extended to operator learning. FiLM was originally introduced for visual reasoning and has since been incorporated into architectures such as graph neural networks [brockschmidt2020gnn]. In the present work, a FiLM-augmented DeepONet surrogate model is employed to reconstruct the full histories of DoC, viscosity, and deformation, despite the fact that only the final deformation at the end of the cure cycle is available experimentally. To address the mismatch between the time discretization used in simulations and that available in experiments, a transfer learning strategy [pan2020transfer] is adopted, in which the DeepONet is first trained on high-fidelity simulation data and subsequently fine-tuned using experimental measurements through a loss function that enforces agreement between the predicted and measured final deformation.
@@
While the FiLM-DeepONet framework provides accurate predictions for the evolution of DoC, viscosity, and deformation, performing reliable uncertainty quantification (UQ) for operator learning remains substantially more challenging. The associated inference problem is typically a very high-dimensional Bayesian inverse problem, which renders gradient-based sampling methods prohibitively expensive. Classical Bayesian approaches such as Hamiltonian Monte Carlo (HMC) [neal2012bayesian] and the Laplace approximation [mackay1992practical] can, in principle, deliver high-quality uncertainty estimates, but they scale poorly, even when diagonal approximations of the Hessian are employed. As a practical alternative, Lakshminarayanan et al. [lakshminarayanan2017simple] proposed deep ensembles trained with proper scoring rules, which can yield well-calibrated predictive distributions with only minor modifications to standard training pipelines. However, these ensemble methods have been criticized for lacking a clear Bayesian interpretation [wilson2022evaluating, reich2013nonparametric, pearce2018uncertainty]. To help bridge this gap, Pearce et al. [pearce2020uncertainty] introduced an ensemble scheme based on approximate Bayesian inference, in which model parameters are regularized around samples from a prior-like distribution, leading to more interpretable uncertainty estimates.
@@
Another promising direction is ensemble Kalman inversion (EKI), a gradient-free ensemble method that has been successfully combined with UQ in scientific machine learning [gao2025scalable, pensoneault2025uncertainty]. Owing to its favorable scaling and rapid convergence, EKI is particularly well suited to high-dimensional operator-learning problems. In this work, we therefore employ EKI to train DeepONet while simultaneously obtaining uncertainty estimates. To further reduce the effective parameter dimension and improve computational efficiency, conventional multilayer perceptrons (MLPs) are replaced with Chebyshev enhanced Kolmogorov Arnold Networks (cKANs) within the DeepONet architecture. The cKANs employed here follow a modified DeepOKAN design, in which Chebyshev polynomials are incorporated into Kolmogorov Arnold Networks (KANs) [liu2024kan, shukla2024comprehensive, faroughi2025scientific] within a standard DeepONet framework. Inspired by Kolmogorov networks [sprecher2002space, koppen2002training], KANs are designed to adapt their activation patterns to the input data, thereby enabling more efficient forward evaluations and, consequently, faster generation of ensemble samples. The model developed as part of this study, thus couples operator learning with transfer learning by training a DeepONet surrogate on simulation data and refining it with experimental measurements to recover full deformation fields from limited observations, while EKI-based training simultaneously yields uncertainty estimates, providing experimental predictions that are both accurate and uncertainty-aware.
@@ 2.1 Experimental Approach
Asymmetric laminates inherently develop residual stresses during cure, resulting in measurable out-of-plane deformation at the end of the process [6_White1993]. To leverage this effect, an asymmetric $[0/90]$ cross-ply laminate configuration was selected. This layup provides a clear and reproducible manifestation of PID, enabling direct comparison between simulations and experiments. The simulation framework used in this study captures the physical mechanisms governing residual stress evolution and the resulting PID. Using this framework, two optimized non-isothermal cure cycles were identified. Experiments were then conducted in which the baseline cure cycle, along with the two optimized cycles, were implemented on a laboratory-scale setup, and the resulting PID was measured. These experiments serve two purposes: (i) to validate the simulation model by confirming that the initial and boundary conditions correctly represent the physical process, and (ii) to provide experimental data that complement the larger simulation-generated dataset for DeepONet model development.
@@ 2.2 Materials and Methods
Two differential scanning calorimetry (DSC) methods are typically used to determine $\alpha_{i}$: (1) measuring the residual heat of reaction in the as-received material, or (2) assessing the shift in the glass transition temperature ($T_{g}$) [19_Sun2002]. In this study, the first method was adopted, wherein the residual heat of reaction was measured to determine $\alpha_{i}$, following the approach described by Sun et al. [19_Sun2002]. Hargis et al. [20_Hargis2006] performed a detailed calorimetric analysis of AS4/3501-6 using DSC to characterize the cure kinetics. Since cure kinetics generally describe the behavior of the thermoset epoxy resin matrix, it is important to accurately determine the resin mass fraction in the prepreg when DSC samples are taken directly from fiber-reinforced prepregs [21_Kim2002]. A more common alternative is to obtain the thermoset epoxy system from the manufacturer as separate, unmixed components, mix them immediately prior to the DSC run, and perform DSC on the neat epoxy system [22_IlLee1982, 23_Chern2002]. However, this latter approach yields the total heat of reaction for freshly mixed resin and may not accurately represent the state of the epoxy in the prepreg, where some prereaction has already taken place. Consequently, several studies have instead performed DSC measurements directly on AS4/3501-6 prepreg and calculated the heat of reaction while compensating for the fiber mass fraction present in the DSC sample [21_Kim2002, 23_Chern2002]. The values reported in the literature are summarized in Table 1.
@@
Lee et al. [22_IlLee1982] 1982 3501-6 Resin @ 20 °C/min to 326.85 °C $473.6\pm 5.4$ NA
Hou et al. [24_Hou1988] 1988 3501-6 Resin @ 20 °C/min $502\pm 21$ NA
White et al. [25_White1993] 1993 AS4 CF/3501-6 Prepreg (42% resin mass) @ 20 °C/min to 350 °C 182.87 (prepreg) 435 (comp. for resin) NA
Kim et al. [21_Kim2002] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 20 °C/min to 350 °C 433.7 (comp. for resin) NA
Kim et al. [21_Kim2002] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 2 °C/min to 350 °C 456.7 (comp. for resin) NA
Chern et al. [23_Chern2002] 2002 3501-6 Resin @ 20 °C/min $508\pm 19$ 2% cure (manufacturer’s note)
Hargis et al. [20_Hargis2006] 2006 3501-6 Resin @ 20 °C/min $382.5\pm 20$ Significant pre-reaction reported
@@
The material used in the compression molding trials was characterized under comparable DSC conditions using a Netzsch Polyma 214 DSC (Netzsch Instruments Inc., Burlington, MA, USA), with nitrogen purge gas and an empty aluminum pan as reference. The results from the first dynamic heating run at 20 °C/min are presented in Figure 1 and show good agreement with the prepreg values reported by White et al. [25_White1993] (see Table 1).
@@ 2.4 Experimental Cure Cycles and Measurements
Figure 4: Temperature–time profiles for the three cure cycles: an isothermal baseline cycle and two non-isothermal cycles identified through the optimization scheme in [18_Limaye2024]. The curves show the measured specimen temperature, the thermal chamber environment, and the corresponding simulated cure profiles.
@@
Table 2: Comparison of experimentally measured and simulated PID (mm) for the three cure cycles: an isothermal baseline cycle and two non-isothermal cycles identified through the optimization scheme in [18_Limaye2024].
@@ 3 Model and Simulations
A comprehensive study by Limaye et al. [18_Limaye2024] demonstrated that the cure cycle governs phase transitions in polymer laminates, progressing from viscous to viscoelastic and ultimately to elastic behavior. The study further highlighted the critical role of thermo-mechanical interactions in modulus evolution and the development of residual stresses and deformations. These coupled phenomena were represented using Polymerization–Gelation–Vitrification (PGV) plots, which characterize material behavior as a function of the applied cure cycle. To capture these complex interactions, a computational cure analysis framework was developed using ABAQUS and COMPRO in a sequential two-step approach, comprising thermo-chemical analysis followed by stress–deformation analysis. The model predicted temperature evolution, DoC, and laminate deformation. These predictions were validated against existing literature, showing strong agreement in both DoC profiles and deformation measurements.
@@ 4 Operator Learning
Operator learning provides a machine learning framework for approximating mappings between function spaces rather than finite-dimensional vectors [lu2019deeponet, lu2024bridging, shih2024transformers, kovachki2023neural, CiCP-35-1194]. The goal is to learn solution operators that map input functions, such as initial or boundary conditions, to output functions, for example the corresponding solutions of partial differential equations. A prototypical architecture for this task is DeepONet [lu2019deeponet], which employs a branch–trunk structure, the branch network encodes the input function, while the trunk network parameterizes the dependence on spatial or temporal query points.

### v3

@@ 1 Introduction
Thermoset fiber reinforced composites are widely adopted for lightweight structural applications in aerospace[57], wind energy [41], automotive[7, 8, 50], prosthetics, wearable electronics and sports equipment. A common route for their manufacturing is the lay-up of pre-impregnated sheets of unidirectional fibers and subsequently curing them under elevated pressure and temperature. During this process, residual stresses are induced due to the mismatch between the thermal expansion coefficients of the fiber and the matrix, as well as the shrinkage of the thermoset matrix as it cures and cross-links, thus affecting the mechanical properties of the manufactured component [62, 28, 66, 2, 14]. Release of some of these residual stresses due to material compliance and demolding results in process induced deformation (PID) that can render the manufactured part out of design tolerances and defective. Additional stresses may also be induced if the deformed and out-of-tolerance part is assembled and fastened in place, potentially leading to premature or catastrophic failure. Accounting for PID and designing the cure cycle for PID reduction is therefore a critical aspect of the design and manufacturing of composite parts. Residual stresses have been shown to reduce by as much as 25 % - 30 % through optimization of the cure cycle [63]. Several studies have tried to develop predictive capabilities using numerical methods. Constitutive laws have been proposed and demonstrated as part of finite element schemes to model cure dependent viscosity development and evolution of modulus and corresponding residual stresses [29, 9, 36, 67]. White et al. [62] investigated, both experimentally and analytically (on LamCure), the effect of changing the manufacturer recommended cure cycle on the resulting curvature for asymmetric cross-ply laminates. Li et al. [30] utilized a FE model for the multi-physics simulation to compute residual stresses, which was then modulated by an improved genetic algorithm to inversely determine the optimal process parameters to minimize residual stresses. Wang et al. [60] performed a fully coupled thermomechanical analysis to model the influence of cure cycles on residual stresses and degree of cure (DOC). They inferred that the temperature and duration of the second dwell had the greatest effect on residual stresses. Hui et al. [17] developed a collaborative multi-objective optimization strategy combining finite element-based cure process analysis with NGSA-II to demonstrate that the temperature gradient, residual stress, and process time can be reduced simultaneously.
@@
Along with advances in physics-based modeling, rapid progress in machine learning (ML) has transformed manufacturing research by enabling efficient data-driven modeling, prediction, and optimization [3, 1, 44, 21]. Deep learning, in particular, has become a central tool across science and engineering. A growing body of work focuses on operator learning, which aims to approximate mappings between input and output functions, a setting that naturally arises in many scientific and engineering applications. For PID in composite structures, a range of ML architectures has been explored. Lin et al. [33] used neural networks, support vector regression, and $k$-nearest neighbors models to rapidly estimate differences in DoC from process parameters, while Fan et al. [11] employed convolutional neural networks to predict PID fields for laminates with varying stacking sequences. Theory-guided and simulation-informed approaches have also been proposed, including theory-guided ML with Gaussian process regression for process optimization [53], random forests for optimal cure design in co-cured composite structures [25], and simulation frameworks that capture competing mechanisms of residual stress evolution [32]. More recently, Liu et al. [35] combined viscoelastic finite element analysis, feature selection (Random Forest, LASSO), and artificial neural networks for rapid PID prediction, while Zhang et al. [65] introduced deep learning models tailored to complex thermo–chemo–mechanical interactions.
@@
Operator learning provides a general ML framework for approximating mappings between function spaces rather than finite-dimensional vectors [38, 39, 54, 23, 34, 59, 4]. A prototypical architecture for this task is the Deep Operator Network (DeepONet) [38], which employs a branch–trunk structure, the branch network encodes the input function, and the trunk network parameterizes the dependence on spatial or temporal query points. Several alternative neural operator architectures have been proposed, including Fourier Neural Operators (FNO) [31] and U-Net-based operator models [52], which mainly differ in how they represent and propagate functional information. Physics-informed variants of DeepONet incorporate PDE residuals and other physical constraints into the loss function, enabling physically consistent predictions in data-scarce regimes [61, 20]. The Two-step DeepONet [26, 20, 49] extends this idea by composing multiple operator stages, thus increasing expressivity for complex multiscale or highly nonlinear mappings while retaining an operator-centric viewpoint.
@@
Building on these developments, Feature-wise Linear Modulation (FiLM) [48] is extended to operator learning. FiLM was originally introduced for visual reasoning and has since been incorporated into architectures such as graph neural networks [5]. In the present work, a FiLM-augmented DeepONet surrogate model is employed to reconstruct the full histories of DoC, viscosity, and deformation, despite the fact that only the final deformation at the end of the cure cycle is available experimentally. To address the mismatch between the time discretization used in simulations and that available in experiments, a transfer learning strategy [43] is adopted, in which the DeepONet is first trained on high-fidelity simulation data and subsequently fine-tuned using experimental measurements through a loss function that enforces agreement between the predicted and measured final deformation.
@@
While the FiLM-DeepONet framework provides accurate predictions for the evolution of DoC, viscosity, and deformation, performing reliable uncertainty quantification (UQ) for operator learning remains substantially more challenging. The associated inference problem is typically a very high-dimensional Bayesian inverse problem, which renders gradient-based sampling methods prohibitively expensive. Classical Bayesian approaches such as Hamiltonian Monte Carlo (HMC) [42] and the Laplace approximation [40] can, in principle, deliver high-quality uncertainty estimates, but they scale poorly, even when diagonal approximations of the Hessian are employed. As a practical alternative, Lakshminarayanan et al. [24] proposed deep ensembles trained with proper scoring rules, which can yield well-calibrated predictive distributions with only minor modifications to standard training pipelines. However, these ensemble methods have been criticized for lacking a clear Bayesian interpretation [64, 51, 46]. To help bridge this gap, Pearce et al. [45] introduced an ensemble scheme based on approximate Bayesian inference, in which model parameters are regularized around samples from a prior-like distribution, leading to more interpretable uncertainty estimates.
@@
Another promising direction is ensemble Kalman inversion (EKI), a gradient-free ensemble method that has been successfully combined with UQ in scientific machine learning [13, 47]. Owing to its favorable scaling and rapid convergence, EKI is particularly well suited to high-dimensional operator-learning problems. In this work, we therefore employ EKI to train DeepONet while simultaneously obtaining uncertainty estimates. To further reduce the effective parameter dimension and improve computational efficiency, conventional multilayer perceptrons (MLPs) are replaced with Chebyshev enhanced Kolmogorov Arnold Networks (cKANs) within the DeepONet architecture. The cKANs employed here follow a modified DeepOKAN design, in which Chebyshev polynomials are incorporated into Kolmogorov Arnold Networks (KANs) [37, 55, 12] within a standard DeepONet framework. Inspired by Kolmogorov networks [56, 22], KANs are designed to adapt their activation patterns to the input data, thereby enabling more efficient forward evaluations and, consequently, faster generation of ensemble samples. The model developed as part of this study, thus couples operator learning with transfer learning by training a DeepONet surrogate on simulation data and refining it with experimental measurements to recover full deformation fields from limited observations, while EKI-based training simultaneously yields uncertainty estimates, providing experimental predictions that are both accurate and uncertainty-aware.
@@ 2.1 Experimental Approach
Asymmetric laminates inherently develop residual stresses during cure, resulting in measurable out-of-plane deformation at the end of the process [62]. To leverage this effect, an asymmetric $[0/90]$ cross-ply laminate configuration was selected. This layup provides a clear and reproducible manifestation of PID, enabling direct comparison between simulations and experiments. The simulation framework used in this study captures the physical mechanisms governing residual stress evolution and the resulting PID. Using this framework, two optimized non-isothermal cure cycles were identified. Experiments were then conducted in which the baseline cure cycle, along with the two optimized cycles, were implemented on a laboratory-scale setup, and the resulting PID was measured. These experiments serve two purposes: (i) to validate the simulation model by confirming that the initial and boundary conditions correctly represent the physical process, and (ii) to provide experimental data that complement the larger simulation-generated dataset for DeepONet model development.
@@ 2.2 Materials and Methods
Two differential scanning calorimetry (DSC) methods are typically used to determine $\alpha_{i}$: (1) measuring the residual heat of reaction in the as-received material, or (2) assessing the shift in the glass transition temperature ($T_{g}$) [58]. In this study, the first method was adopted, wherein the residual heat of reaction was measured to determine $\alpha_{i}$, following the approach described by Sun et al. [58]. Hargis et al. [15] performed a detailed calorimetric analysis of AS4/3501-6 using DSC to characterize the cure kinetics. Since cure kinetics generally describe the behavior of the thermoset epoxy resin matrix, it is important to accurately determine the resin mass fraction in the prepreg when DSC samples are taken directly from fiber-reinforced prepregs [19]. A more common alternative is to obtain the thermoset epoxy system from the manufacturer as separate, unmixed components, mix them immediately prior to the DSC run, and perform DSC on the neat epoxy system [27, 6]. However, this latter approach yields the total heat of reaction for freshly mixed resin and may not accurately represent the state of the epoxy in the prepreg, where some prereaction has already taken place. Consequently, several studies have instead performed DSC measurements directly on AS4/3501-6 prepreg and calculated the heat of reaction while compensating for the fiber mass fraction present in the DSC sample [19, 6]. The values reported in the literature are summarized in Table 1.
@@
Lee et al. [27] 1982 3501-6 Resin @ 20 °C/min to 326.85 °C $473.6\pm 5.4$ NA
Hou et al. [16] 1988 3501-6 Resin @ 20 °C/min $502\pm 21$ NA
White et al. [63] 1993 AS4 CF/3501-6 Prepreg (42% resin mass) @ 20 °C/min to 350 °C 182.87 (prepreg) 435 (comp. for resin) NA
Kim et al. [19] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 20 °C/min to 350 °C 433.7 (comp. for resin) NA
Kim et al. [19] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 2 °C/min to 350 °C 456.7 (comp. for resin) NA
Chern et al. [6] 2002 3501-6 Resin @ 20 °C/min $508\pm 19$ 2% cure (manufacturer’s note)
Hargis et al. [15] 2006 3501-6 Resin @ 20 °C/min $382.5\pm 20$ Significant pre-reaction reported
@@
The material used in the compression molding trials was characterized under comparable DSC conditions using a Netzsch Polyma 214 DSC (Netzsch Instruments Inc., Burlington, MA, USA), with nitrogen purge gas and an empty aluminum pan as reference. The results from the first dynamic heating run at 20 °C/min are presented in Figure 1 and show good agreement with the prepreg values reported by White et al. [63] (see Table 1).
@@ 2.4 Experimental Cure Cycles and Measurements
Figure 4: Temperature–time profiles for the three cure cycles: an isothermal baseline cycle and two non-isothermal cycles identified through the optimization scheme in [32]. The curves show the measured specimen temperature, the thermal chamber environment, and the corresponding simulated cure profiles.
@@
Table 2: Comparison of experimentally measured and simulated PID (mm) for the three cure cycles: an isothermal baseline cycle and two non-isothermal cycles identified through the optimization scheme in [32].
@@ 3 Model and Simulations
A comprehensive study by Limaye et al. [32] demonstrated that the cure cycle governs phase transitions in polymer laminates, progressing from viscous to viscoelastic and ultimately to elastic behavior. The study further highlighted the critical role of thermo-mechanical interactions in modulus evolution and the development of residual stresses and deformations. These coupled phenomena were represented using Polymerization–Gelation–Vitrification (PGV) plots, which characterize material behavior as a function of the applied cure cycle. To capture these complex interactions, a computational cure analysis framework was developed using ABAQUS and COMPRO in a sequential two-step approach, comprising thermo-chemical analysis followed by stress–deformation analysis. The model predicted temperature evolution, DoC, and laminate deformation. These predictions were validated against existing literature, showing strong agreement in both DoC profiles and deformation measurements.
@@ 4 Operator Learning
Operator learning provides a machine learning framework for approximating mappings between function spaces rather than finite-dimensional vectors [38, 39, 54, 23, 34]. The goal is to learn solution operators that map input functions, such as initial or boundary conditions, to output functions, for example the corresponding solutions of partial differential equations. A prototypical architecture for this task is DeepONet [38], which employs a branch–trunk structure, the branch network encodes the input function, while the trunk network parameterizes the dependence on spatial or temporal query points.